# PointAD+: Learning Hierarchical Representations for ZS 3D Anomaly Detection

### v4

@@ Abstract
Zero-shot (ZS) 3D anomaly detection is a critical yet underexplored field that addresses scenarios where target 3D training samples are unavailable. This presents a challenge, as it requires the model to accurately detect 3D anomalies in previously unseen categories. In this paper, we aim to transfer CLIP’s robust 2D generalization capabilities to identify 3D anomalies across unseen objects of highly diverse class semantics. We first design PointAD, which leverages point-pixel correspondence to represent 3D anomalies through their associated rendering pixel representations. This representation is referred to as implicit point representation, as it focuses solely on rendering pixel anomalies but neglects the inherent spatial relationships within point clouds. However, spatial awareness modeling is important for accurate 3D abnormality recognition. Therefore, we propose PointAD+ to further broaden the interpretation of 3D anomalies by introducing explicit point representation that is aggregated from implicit point representation, modeling abnormal point relative position. We further introduce G-aggregation to involve geometry information to enable the explicit point representation to be spatially aware. To simultaneously capture rendering and geometry abnormality, PointAD+ proposes hierarchical representation learning, incorporating implicit and explicit anomaly semantics into hierarchical text prompts: rendering prompts for the rendering layer and geometry prompts for the geometry layer. A cross-hierarchy contrastive alignment is introduced to promote the interaction between the rendering and geometry layers, facilitating mutual anomaly learning. Finally, PointAD+ integrates anomaly semantics from both layers to capture the generalized anomaly semantics. During the test, PointAD+ can integrate RGB information in a plug-and-play manner and further improve its detection performance. Extensive experiments demonstrate the superiority of PointAD+ in ZS 3D anomaly detection across unseen objects with highly diverse class semantics, achieving a holistic understanding of abnormality.
@@ 1 Introduction
Anomaly detection has garnered considerable attention from both academia and industry. It has been widely applied to diverse applications, such as industrial inspection [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15] and medical analysis [16, 17, 18]. Driven by the growing need to understand 3D objects in real-world applications, such as spatial intelligence, 3D anomaly detection has emerged as a promising field [19, 20, 21, 22, 23, 24, 25, 26, 27]. Unlike conventional 2D anomaly detection methods that rely solely on RGB information [19, 20, 21, 22, 23, 25, 28, 29, 30], 3D anomaly detection focuses on uncovering spatial relationships that reveal abnormal patterns. This approach is particularly effective when normal regions share similar RGB characteristics with anomalies or when defects closely mimic the appearance of an object’s foreground. For instance, as shown in Figure 1, chocolates (in black) on cookies closely resemble hole anomalies (also in black), causing the RGB model to mistakenly classify the chocolates as anomalies. Similarly, surface damage on a potato exhibits a color pattern similar to that of the potato’s foreground, leading to misdetection by the RGB model. In contrast, 3D anomaly detection can effectively mitigate such failures by identifying abnormal spatial relationships.
@@
Current 3D anomaly detection methods typically operate under an unsupervised setting, assuming that the target point clouds are available and entirely normal. To effectively model the normal decision boundary, these methods generally store normal point features from the training dataset as a reference memory and identify anomalies by measuring the distance between the test features and the stored ones [19, 23, 25]. However, this assumption does not hold in many situations where training samples in the target dataset are inaccessible due to privacy protection (e.g., involvement of trade secrets) or the absence of target training data (e.g., a new product that has never been seen before) [16]. The unsupervised paradigm detects class-level anomalies by memorizing or reconstructing normal point features of the target class. Hence, it is limited in its ability to capture anomalies in unseen objects. Noteably, ZS 3D anomaly detection requires the model to identify anomalies in unseen classes. While ZS anomaly detection has been explored in 2D RGB images [31, 16, 32], ZS 3D anomaly detection remains an underexplored research area. This task is particularly challenging as it requires the model to capture generic anomaly semantics from a limited set of auxiliary classes, despite significant discrepancies in object semantics and the presence of unknown anomaly patterns. Recently, Vision-Language Models (VLMs) have shown strong generalization capabilities to unseen objects and have been successfully applied to various downstream tasks [33, 34, 35, 36, 37, 38]. One popular model, CLIP, has demonstrated impressive ZS performance in detecting 2D anomalies [31, 11, 16]. Integrating CLIP into the detection model presents a potential solution to the challenging yet unexplored domain of ZS 3D anomaly detection. It is nontrivial to directly leverage CLIP for ZS 3D anomaly detection as the knowledge gap between the 2D domain and 3D domain is significant.
@@
PointAD relies on view-wise 2D features to represent points indirectly. However, CLIP itself cannot capture the geometric information embedded in point clouds. Such spatial awareness is crucial for modeling accurate point representations, particularly for reliable abnormality recognition. PointAD+ instead characterizes 3D abnormality in two aspects: implicit 3D abnormality on the rendering layer and explicit 3D abnormality on the geometry layer. Holding these two aspects of 3D anomalies, we propose PointAD+ that learns hierarchical representations: (1) align implicit point representations on the rendering layer, which interpret each point through aggregated pixel-level logits from multi-view renderings; and (2) align explicit point representations on the geometry layer, aggregated from implicit point representations and explicitly encoding spatial geometry.
@@ 2 Related Work
3D anomaly detection has recently gained attention in the community of anomaly detection. Unlike 2D anomaly detection, which leverages RGB information, 3D anomaly detection focuses on identifying unusual spatial relationships in 3D objects. While 2D anomaly detection has been extensively explored, research on 3D anomaly detection remains relatively limited. This disparity is largely due to the challenges associated with acquiring high-quality 3D data, as collecting 3D datasets is more costly and labor-intensive compared to 2D images. MVTec 3D-AD [20] is the first public dataset, which provides the point cloud anomalies and the corresponding 2D-RGB information. Subsequently, Eyecandies [39] and Real3D-AD [40] provide the 3D dataset for a wide exploration [41]. These datasets bridge the connection between 3D and 2D anomaly detection. 3D-ST [20] uses a teacher net to extract dense local geometry descriptors and design a student net to match such descriptors. AST [21] introduces asymmetric teacher and student net to further improve 3D anomaly detection. Instead of only using point clouds, BTF [19], multimodal 3DM [23, 42], CPFM [43], and SDM [25] integrate point features and RGB pixel features to detect 3D anomalies. While these approaches exhibit commendable performance by storing object-specific normal point and pixel features within the unsupervised learning framework, such paradigms simultaneously limit their generalization capacity to point clouds from unseen objects, which is crucial to detecting anomalies when the target object is unavailable. To fill this gap, we introduce PointAD+ and PointAD to identify unseen anomalies across diverse objects. They extend CLIP to the realm of ZS 3D and show robust generalization in capturing generic normality and abnormality within points. Furthermore, PointAD+ and PointAD serve as a unified framework, allowing seamless integration of 3D and RGB without additional training.
@@
The approaches for representing 3D objects can be broadly categorized into two branches. One way directly employs point-based networks to model 3D objects. Methods like PointNet [44] and PointNet++ [45] model the relationships between individual points to extract 3D features. The other way involves converting 3D data into a 2D format [46, 47] by rendering point clouds into multiple views. This allows the use of robust 2D backbones to process the 2D renderings and subsequently handle 3D information. PointCLIP [48] directly projects raw points onto image planes for efficiency, however, this method often results in depth maps that lack detailed geometry information. Conversely, rendering-based methods [46, 49] generate high-quality 2D renderings from point clouds, preserving detailed local semantics. More related work in unsupervised 3D anomaly detection includes 3DSR [50], which extracts depth images as 3D information, and CPFM [43], which utilizes normal features from 2D renderings. In this paper, we adopt a rendering strategy for source point clouds and their annotations to obtain high-quality 2D rendering information and corresponding supervision. This enables PointAD+ and PointAD to capture generic anomaly semantics better, facilitating the recognition of abnormality in unseen objects.
@@ 2.3 Prompt Learning
With the increase in model parameters, training a model from scratch requires substantial computational resources and time. This high training cost hinders the rapid development of deep learning technologies. Therefore, rather than training or fine-tuning the entire network, prompt learning focuses on optimizing a minimal set of parameters to adapt the network for downstream tasks. CoOp, a seminal work in prompt learning [34, 51], introduces global context optimization to update learnable text prompts for few-shot recognition. DenseCLIP [35] extends this approach to dense classifications. More recently, AnomalyCLIP [16] introduces object-agnostic prompt learning to capture generic normality and abnormality in images for 2D anomaly detection. Other works, such as [52] and [53], have built upon this by incorporating additional vision information. In contrast, PointAD+ innovatively transforms the flat prompts used in these methods into hierarchical prompts to capture both rendering and spatial abnormality for ZS 3D anomaly detection. Leveraging this hierarchical structure, we introduce hierarchy representation learning to achieve an in-depth understanding of ZS 3D and multimodal 3D anomaly detection.
@@ 3.1 A Review of CLIP and AnomalyCLIP
Replacing fixed text prompts with learnable text prompts is an effective method to adapt CLIP for downstream tasks. AnomalyCLIP [16] introduces object-agnostic text prompts to adapt CLIP for ZS 2D anomaly detection. The success of AnomalyCLIP demonstrates the promise of adapting CLIP for ZSAD by utilizing a lightweight, learnable text prompt. However, ZS 3D anomaly detection remains underexplored, primarily because there is no 3D foundation model with the strong generalization capabilities of CLIP. In this paper, we explore how to transfer CLIP’s strong generalization ability to 3D anomaly detection in a ZS manner.
@@ 3.3 Motivation of PointAD+
PointAD employs implicit point representations to capture 3D anomaly semantics by aggregating pixel-level abnormalities into point-level ones. However, when anomalies are not well reflected in renderings, this approach can lead to missed anomalies and false alarms (Figure LABEL:fig:_ei1). In contrast, explicit point segmentation encodes geometric information to localize anomaly regions with higher precision, as it is more sensitive to spatial variations. Yet, explicit point learning is not universally superior: as shown in Figure LABEL:fig:_ei2, it struggles with edge points, where abrupt spatial changes often trigger false alarms. Implicit point representations, by relying on 2D features, are less sensitive to such edge variations and avoid this problem. To exploit the strengths of both, PointAD+ introduces hierarchical representation learning that integrates implicit point learning on the rendering layer and explicit point learning on the geometry layer, enabling a more comprehensive understanding of abnormalities.
@@ 3.4 Multi-view Rendering
Multi-view projection is a crucial technique for understanding point clouds from 2D perspectives. Some multi-view projection approaches project point clouds into various depth maps, providing adequate shape information for class recognition [48]. However, in this paper, the objective of PointAD+ and PointAD is to learn both generic global and local anomaly semantics. Depth-map projection lacks sufficient resolution to represent fine-grained anomaly semantics accurately. Hence, we adopt high-precision rendering to preserve the original 3D information offline. Specifically, given an auxiliary dataset of point clouds $\mathcal{D}_{3d}=\{(x^{3d}_{i},y^{3d}_{i})\}_{i=1}^{N}$, we define the rendering matrix as $R^{(k)}$ for the $k$-thth view, with a total of $K$ views. We simultaneously render point clouds and point-level ground truths from different views to obtain their corresponding 2D renderings, which is given by $x_{i}^{(k)}=R^{(k)}(x^{3d}_{i})$ and $y_{i}^{(k)}=R^{(k)}(y^{3d}_{i})$, where $x_{i}^{(k)}\in\mathbb{R}^{H\times W}$ and $y_{i}^{(k)}\in\mathbb{R}^{H\times W}$ respectively represent the $k$-thth 2D rendering and corresponding pixel-level ground truth in the $i$-thth point cloud, i.e., 0 for normality and 1 for abnormality.
@@ 3.6 Hierarchical representation learning
Specifically, we design two types of learnable text prompts: rendering prompts $g^{r}_{c}$ and geometry prompts $g^{g}_{c}$, each comprising both normality and abnormality components. Rendering prompts are designed to learn implicit point anomalies by leveraging point-pixel correspondence to represent points through their associated rendering pixel representations. In contrast, geometry prompts capture abnormal spatial relationships by aligning aggregated point representations enriched with geometric information within point clouds. PointAD initializes these prompts in AnomalyCLIP [16] manner:

### v5

@@ Abstract
Zero-shot (ZS) 3D anomaly detection is a critical yet underexplored field that addresses scenarios where target 3D training samples are unavailable. This presents a challenge, as it requires the model to accurately detect 3D anomalies in previously unseen categories. In this paper, we aim to transfer CLIP’s robust 2D generalization capabilities to identify 3D anomalies across unseen objects of highly diverse class semantics. We first design PointAD, which leverages point-pixel correspondence to represent 3D anomalies through their associated rendering pixel representations. This representation is referred to as implicit point representation, as it focuses solely on rendering pixel anomalies but neglects the inherent spatial relationships within point clouds. However, spatial awareness modeling is important for accurate 3D abnormality recognition. Therefore, we propose PointAD+ to further broaden the interpretation of 3D anomalies by introducing explicit point representation that is aggregated from implicit point representation, modeling abnormal point relative position. We introduce G-aggregation to involve geometry information to enable the explicit point representation to be spatially aware. To simultaneously capture rendering and geometry abnormality, PointAD+ proposes hierarchical representation learning, incorporating implicit and explicit anomaly semantics into hierarchical text prompts: rendering prompts for the rendering layer and geometry prompts for the geometry layer. A cross-hierarchy contrastive alignment is introduced to promote the interaction between the rendering and geometry layers, facilitating mutual anomaly learning. Finally, PointAD+ integrates anomaly semantics from both layers to capture the generalized anomaly semantics. During the test, PointAD+ can integrate RGB information in a plug-and-play manner and further improve its detection performance. Extensive experiments demonstrate the superiority of PointAD+ in ZS 3D anomaly detection across unseen objects with highly diverse class semantics, achieving a holistic understanding of abnormality.
@@ 1 Introduction
Anomaly detection has garnered considerable attention from both academia and industry. It has been widely applied to diverse applications, such as industrial inspection [bergmann2019mvtec, bergmann2020uninformed, liznerski2020explainable, pang2021explainable, Zhou2023PullP, li2024clipsam, pang2021deep, roth2022towards, zuo2024clip3d, deng2022anomaly, jeong2023winclip, xie2023pushing, cao2023anomaly, gu2023anomalygpt, zhou2024pointad] and medical analysis [zhou2023anomalyclip, tian2023self, tian2021constrained]. Driven by the growing need to understand 3D objects in real-world applications, such as spatial intelligence, 3D anomaly detection has emerged as a promising field [horwitz2023back, bergmann2023anomaly, rudolph2023asymmetric, chen2023easynet, wang2023multimoda, cheng2024zeroshotpointcloudanomaly, chu2023shape, ye2025po3ad, liang2025examining]. Unlike conventional 2D anomaly detection methods that rely solely on RGB information [horwitz2023back, bergmann2023anomaly, rudolph2023asymmetric, chen2023easynet, wang2023multimoda, chu2023shape, huang2022registration, you2022unified, sun2023when], 3D anomaly detection focuses on uncovering spatial relationships that reveal abnormal patterns. This approach is particularly effective when normal regions share similar RGB characteristics with anomalies or when defects closely mimic the appearance of an object’s foreground. For instance, as shown in Figure 1, chocolates (in black) on cookies closely resemble hole anomalies (also in black), causing the RGB model to mistakenly classify the chocolates as anomalies. Similarly, surface damage on a potato exhibits a color pattern similar to that of the potato’s foreground, leading to misdetection by the RGB model. In contrast, 3D anomaly detection can effectively mitigate such failures by identifying abnormal spatial relationships.
@@
Current 3D anomaly detection methods typically operate under an unsupervised setting, assuming that the target point clouds are available and entirely normal. To effectively model the normal decision boundary, these methods generally store normal point features from the training dataset as a reference memory and identify anomalies by measuring the distance between the test features and the stored ones [horwitz2023back, wang2023multimoda, chu2023shape]. However, this assumption does not hold in many situations where training samples in the target dataset are inaccessible due to privacy protection (e.g., involvement of trade secrets) or the absence of target training data (e.g., a new product that has never been seen before) [zhou2023anomalyclip]. The unsupervised paradigm detects class-level anomalies by memorizing or reconstructing normal point features of the target class. Hence, it is limited in its ability to capture anomalies in unseen objects. Noteably, ZS 3D anomaly detection requires the model to identify anomalies in unseen classes. While ZS anomaly detection has been explored in 2D RGB images [ming2022delving, zhou2023anomalyclip, li2024promptad], ZS 3D anomaly detection remains an underexplored research area. This task is particularly challenging as it requires the model to capture generic anomaly semantics from a limited set of auxiliary classes, despite significant discrepancies in object semantics and the presence of unknown anomaly patterns. Recently, Vision-Language Models (VLMs) have shown strong generalization capabilities to unseen objects and have been successfully applied to various downstream tasks [radford2021learning, zhou2022learning, rao2022denseclip, sun2022dualcoop, khattak2023maple, kirillov2023segment]. One popular model, CLIP, has demonstrated impressive ZS performance in detecting 2D anomalies [ming2022delving, jeong2023winclip, zhou2023anomalyclip]. Integrating CLIP into the detection model presents a potential solution to the challenging yet unexplored domain of ZS 3D anomaly detection. It is nontrivial to directly leverage CLIP for ZS 3D anomaly detection as the knowledge gap between the 2D domain and 3D domain is significant.
@@
PointAD relies on view-wise 2D renderings to represent points. When point anomalies are not clearly reflected in the rendered views, the implicit point representations struggle to capture them. Such anomalies require to be capable of underlying spatial relations, yet CLIP is inherently limited in modeling the geometric structures present in point clouds. To overcome this limitation, we introduce explicit point representations that directly encode geometric information. Further, we propose PointAD+ that learns hierarchical representations: (1) align implicit point representations on the rendering layer, which interpret each point through aggregated pixel-level logits from multi-view renderings; and (2) align explicit point representations on the geometry layer, aggregated from implicit point representations and explicitly encoding spatial geometry.
@@ 2 Related Work
3D anomaly detection has recently gained attention in the community of anomaly detection. Unlike 2D anomaly detection, which leverages RGB information, 3D anomaly detection focuses on identifying unusual spatial relationships in 3D objects. While 2D anomaly detection has been extensively explored, research on 3D anomaly detection remains relatively limited. This disparity is largely due to the challenges associated with acquiring high-quality 3D data, as collecting 3D datasets is more costly and labor-intensive compared to 2D images. MVTec 3D-AD [bergmann2023anomaly] is the first public dataset, which provides the point cloud anomalies and the corresponding 2D-RGB information. Subsequently, Eyecandies [Bonfiglioli_2022_ACCV] and Real3D-AD [liu2023realdad] provide the 3D dataset for a wide exploration [li2024towards]. These datasets bridge the connection between 3D and 2D anomaly detection. 3D-ST [bergmann2023anomaly] uses a teacher net to extract dense local geometry descriptors and design a student net to match such descriptors. AST [rudolph2023asymmetric] introduces asymmetric teacher and student net to further improve 3D anomaly detection. Instead of only using point clouds, BTF [horwitz2023back], M3DM [wang2023multimoda, wang2025m3dm], CPFM [cao2023complementary], and SDM [chu2023shape] integrate point features and RGB pixel features to detect 3D anomalies. While these approaches exhibit commendable performance by storing object-specific normal point and pixel features within the unsupervised learning framework, such paradigms simultaneously limit their generalization capacity to point clouds from unseen objects, which is crucial to detecting anomalies when the target object is unavailable. To fill this gap, we introduce PointAD+ and PointAD to identify unseen anomalies across diverse objects. They extend CLIP to the realm of ZS 3D and show robust generalization in capturing generic normality and abnormality within points. Furthermore, PointAD+ and PointAD serve as a unified framework, allowing seamless integration of 3D and RGB without additional training.
@@
The approaches for representing 3D objects can be broadly categorized into two branches. One way directly employs point-based networks to model 3D objects. Methods like PointNet [qi2017pointnet] and PointNet++ [qi2017pointnet++] model the relationships between individual points to extract 3D features. The other way involves converting 3D data into a 2D format [su2015multi, goyal2021revisiting] by rendering point clouds into multiple views. This allows the use of robust 2D backbones to process the 2D renderings and subsequently handle 3D information. PointCLIP [zhang2022pointclip] directly projects raw points onto image planes for efficiency, however, this method often results in depth maps that lack detailed geometry information. Conversely, rendering-based methods [su2015multi, hamdi2021mvtn] generate high-quality 2D renderings from point clouds, preserving detailed local semantics. More related work in unsupervised 3D anomaly detection includes 3DSR [zavrtanik2024cheating], which extracts depth images as 3D information, and CPFM [cao2023complementary], which utilizes normal features from 2D renderings. In this paper, we adopt a rendering strategy for source point clouds and their annotations to obtain high-quality 2D rendering information and corresponding supervision. This enables PointAD+ and PointAD to capture generic anomaly semantics better, facilitating the recognition of abnormality in unseen objects.
@@ 2.3 Prompt Learning
With the increase in model parameters, training a model from scratch requires substantial computational resources and time. This high training cost hinders the rapid development of deep learning technologies. Therefore, rather than training or fine-tuning the entire network, prompt learning focuses on optimizing a minimal set of parameters to adapt the network for downstream tasks. CoOp, a seminal work in prompt learning [zhou2022learning, zhou2022conditional], introduces global context optimization to update learnable text prompts for few-shot recognition. DenseCLIP [rao2022denseclip] extends this approach to dense classifications. More recently, AnomalyCLIP [zhou2023anomalyclip] introduces object-agnostic prompt learning to capture generic normality and abnormality in images for 2D anomaly detection. Other works, such as [cao2024adaclip] and [qu2024vcp], have built upon this by incorporating additional vision information. In contrast, PointAD+ innovatively transforms the flat prompts used in these methods into hierarchical prompts to capture both rendering and spatial abnormality for ZS 3D anomaly detection. Leveraging this hierarchical structure, we introduce hierarchy representation learning to achieve an in-depth understanding of ZS 3D and multimodal 3D anomaly detection.
@@ 3.1 A Review of CLIP and AnomalyCLIP
Replacing fixed text prompts with learnable text prompts is an effective method to adapt CLIP for downstream tasks. AnomalyCLIP [zhou2023anomalyclip] introduces object-agnostic text prompts to adapt CLIP for ZS 2D anomaly detection. The success of AnomalyCLIP demonstrates the promise of adapting CLIP for ZSAD by utilizing a lightweight, learnable text prompt. However, ZS 3D anomaly detection remains underexplored, primarily because there is no 3D foundation model with the strong generalization capabilities of CLIP. In this paper, we explore how to transfer CLIP’s strong generalization ability to 3D anomaly detection in a ZS manner.
@@ 3.3 Motivation of PointAD+
PointAD employs implicit point representations to capture 3D anomaly semantics by aggregating pixel-level abnormalities into point-level ones. However, when anomalies are not well reflected in renderings, this approach can lead to missed anomalies and false alarms (Figure LABEL:fig:_ei1). In contrast, explicit point segmentation encodes geometric information to localize anomaly regions with higher precision, as it is more sensitive to spatial variations. Yet, explicit point learning is not universally superior: as shown in Figure LABEL:fig:_ei2, it struggles with edge points, where abrupt spatial changes often trigger false alarms. Implicit point representations, by relying on 2D features, are less sensitive to such edge variations and avoid this problem. To exploit the strengths of both, PointAD+ introduces hierarchical representation learning that integrates implicit point learning on the rendering layer and explicit point learning on the geometry layer, enabling a more comprehensive understanding of abnormalities. Figure LABEL:fig:_ei3 shows that PointAD+ achieves a larger performance gain on Real3D-AD (stereo point clouds) than on MVTec3D-AD and Eyecandies. Real3D-AD contains stereo point clouds, which require the model to be more sensitive to point relations. The significant improvement demonstrates the superiority of PointAD+ in spatial relation modeling.
@@ 3.4 Multi-view Rendering
Multi-view projection is a crucial technique for understanding point clouds from 2D perspectives. Some multi-view projection approaches project point clouds into various depth maps, providing adequate shape information for class recognition [zhang2022pointclip]. However, in this paper, the objective of PointAD+ and PointAD is to learn both generic global and local anomaly semantics. Depth-map projection lacks sufficient resolution to represent fine-grained anomaly semantics accurately. Hence, we adopt high-precision rendering to preserve the original 3D information offline. Specifically, given an auxiliary dataset of point clouds $\mathcal{D}_{3d}=\{(x^{3d}_{i},y^{3d}_{i})\}_{i=1}^{N}$, we define the rendering matrix as $R^{(k)}$ for the $k$-thth view, with a total of $K$ views. We simultaneously render point clouds and point-level ground truths from different views to obtain their corresponding 2D renderings, which is given by $x_{i}^{(k)}=R^{(k)}(x^{3d}_{i})$ and $y_{i}^{(k)}=R^{(k)}(y^{3d}_{i})$, where $x_{i}^{(k)}\in\mathbb{R}^{H\times W}$ and $y_{i}^{(k)}\in\mathbb{R}^{H\times W}$ respectively represent the $k$-thth 2D rendering and corresponding pixel-level ground truth in the $i$-thth point cloud, i.e., 0 for normality and 1 for abnormality.
@@ 3.6 Hierarchical representation learning
Specifically, we design two types of learnable text prompts: rendering prompts $g^{r}_{c}$ and geometry prompts $g^{g}_{c}$, each comprising both normality and abnormality components. Rendering prompts are designed to learn implicit point anomalies by leveraging point-pixel correspondence to represent points through their associated rendering pixel representations. In contrast, geometry prompts capture abnormal spatial relationships by aligning aggregated point representations enriched with geometric information within point clouds. PointAD initializes these prompts in AnomalyCLIP [zhou2023anomalyclip] manner: